OpenStudy (priyar):

The area of the region above the x-axis bounded by the curve y=tanx, {x lies bet. 0 and pi/2} and the tangent to the curve at x=pi/4 is..

OpenStudy (michele_laino):

we have to write the equation of the tangent line, first

OpenStudy (priyar):

ya i got y=2x+1-pi/2

OpenStudy (michele_laino):

correct!

OpenStudy (priyar):

thanks.. my doubt lies in the graph part..

OpenStudy (priyar):

i am not able to find the correct region..

OpenStudy (priyar):

i can do the integration part..if i knew that..

OpenStudy (michele_laino):

now, the x-intercept, is: \[\Large x = \frac{\pi }{4} - \frac{1}{2}\] so we can make this drawing: 

OpenStudy (michele_laino):

I know it is not a good graph

OpenStudy (michele_laino):